### 08 Mar 2024 v1

@@ Abstract
This paper considers image-based virtual try-on, which renders an image of a person wearing a curated garment, given a pair of images depicting the person and the garment, respectively. Previous works adapt existing exemplar-based inpainting diffusion models for virtual try-on to improve the naturalness of the generated visuals compared to other methods (e.g., GAN-based), but they fail to preserve the identity of the garments. To overcome this limitation, we propose a novel diffusion model that improves garment fidelity and generates authentic virtual try-on images. Our method, coined IDM–VTON, uses two different modules to encode the semantics of garment image; given the base UNet of the diffusion model, 1) the high-level semantics extracted from a visual encoder are fused to the cross-attention layer, and then 2) the low-level features extracted from parallel UNet are fused to the self-attention layer. In addition, we provide detailed textual prompts for both garment and person images to enhance the authenticity of the generated visuals. Finally, we present a customization method using a pair of person-garment images, which significantly improves fidelity and authenticity. Our experimental results show that our method outperforms previous approaches (both diffusion-based and GAN-based) in preserving garment details and generating authentic virtual try-on images, both qualitatively and quantitatively. Furthermore, the proposed customization method demonstrates its effectiveness in a real-world scenario.
@@ 4.3 Results on In-the-Wild Dataset
Fig. 5 shows the qualitative results of IDM–VTON (ours) compared to other baselines [^1] We do not compare with GP-VTON as it uses private parsing modules. Comparison with HR-VTON is in the appendix.. We see that IDM–VTON generates more authentic images than LaDI-VTON, DCI-VTON, and Stable VITON, yet it struggles to preserve the intricate patterns of garments. By customizing IDM–VTON, we obtain virtual try-on images with a high degree of consistency to the garment (e.g., the logos and the text renderings), but we find the customization does not work well for StableVITON.

### 19 Mar 2024 v2

@@ Abstract
This paper considers image-based virtual try-on, which renders an image of a person wearing a curated garment, given a pair of images depicting the person and the garment, respectively. Previous works adapt existing exemplar-based inpainting diffusion models for virtual try-on to improve the naturalness of the generated visuals compared to other methods (e.g., GAN-based), but they fail to preserve the identity of the garments. To overcome this limitation, we propose a novel diffusion model that improves garment fidelity and generates authentic virtual try-on images. Our method, coined IDM–VTON, uses two different modules to encode the semantics of garment image; given the base UNet of the diffusion model, 1) the high-level semantics extracted from a visual encoder are fused to the cross-attention layer, and then 2) the low-level features extracted from parallel UNet are fused to the self-attention layer. In addition, we provide detailed textual prompts for both garment and person images to enhance the authenticity of the generated visuals. Finally, we present a customization method using a pair of person-garment images, which significantly improves fidelity and authenticity. Our experimental results show that our method outperforms previous approaches (both diffusion-based and GAN-based) in preserving garment details and generating authentic virtual try-on images, both qualitatively and quantitatively. Furthermore, the proposed customization method demonstrates its effectiveness in a real-world scenario. More visualizations are available in our project page.
@@ 4.3 Results on In-the-Wild Dataset
Fig. 5 shows the qualitative results of IDM–VTON (ours) compared to other baselines [^1] We do not compare with GP-VTON as it uses private parsing modules. Comparison with HR-VITON is in the appendix.. We see that IDM–VTON generates more authentic images than LaDI-VTON, DCI-VTON, and Stable VITON, yet it struggles to preserve the intricate patterns of garments. By customizing IDM–VTON, we obtain virtual try-on images with a high degree of consistency to the garment (e.g., the logos and the text renderings), but we find the customization does not work well for StableVITON.
@@ 5 Conclusion
Appendix
Appendix 0.A Implementation Details
0.A.1 In-the-Wild Dataset
Figure 8: Examples of In-the-Wild dataset. We collect pairs of garment and human wearing the garment.
In-the-Wild dataset comprises multiple human images wearing each target garment. Images of garment are collected from MLB online shopping mall [^2] https://www.mlb-korea.com/main/mall/view and images of human wearing each garment are gathered from social media platforms like Instagram. As shown in Fig. 8, the human images exhibit diverse backgrounds, ranging from parks and buildings to snowy landscapes. For preprocessing, we employ center cropping on the human images to achieve resolutions of $1024\times 768$ resolutions, while the garment images are resized to the same dimensions for compatible setting.
0.A.2 Training and Inference
We train the model using the Adam [23] optimizer with a fixed learning rate of 1e-5 over 130 epochs (63k iterations with batch size of 24). It takes around 40 hours in training with 4$\times$×A800 GPUs. We apply data augmentations following StableVITON [22], where we apply horizontal flip (with probability $0.5$), random affine shifting and scaling (limit of $0.2$, with probability $0.5$) to the inputs of TryonNet, i.e., $\mathbf{x}_{p},\mathbf{x}_{\textrm{pose}},\mathbf{x}_{m}$ and $\mathbf{m}$. For customization, we fine-tune our model using the Adam optimizer with a fixed learning rate of 1e-6 for 100 steps. It takes around 2 minutes with a single A800 GPU.
During the inference, we generate images using the DDPM scheduler with 30 steps. We set the strength value to 1.0, i.e., denoising begins from random Gaussian noise, to ignore the masked portion of the input image. For classifier-free guidance [14] (CFG), we merge both conditioning, i.e., low-level features $\mathbf{L}_{g}$ from GarmentNet and high-level semantics $\mathbf{H}_{g}$ from IP-Adapter, as these conditions contain features of the same garment image, following SpaText [1]. In specific, the forward is given as follows:
$\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t};\mathbf{L}_{g},\mathbf{H}_{g},t)=s\cdot(\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t};\mathbf{L}_{g},\mathbf{H}_{g},t)-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t};t))+\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t};t)\text{,}$ (4)
where $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t};\mathbf{L}_{g},\mathbf{H}_{g},t)$ denotes noise prediction with conditions $\mathbf{L}_{g}$ and $\mathbf{H}_{g}$, and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t};t)$ denotes unconditional noise prediction. We use guidance scale $s$=2.0, which works well in practice.
0.A.3 Detailed captioning of garments
Figure 9: Illustration of generating detailed captions of garments. We utilize pretrained fashion attribute tagging annotator to extract information of garment and generate detailed captions of the garment based on extracted information.
We generate detailed captions for each garment to leverage the prior knowledge of T2I diffusion models. We employ OMNIOUS.AI’s commercial  [^3] https://omnicommerce.ai fashion attribute tagging annotator, which has been trained with over 1,000 different fashion attributes. The image annotator provides various feature categories present in a given image, such as sleeve length and neckline type. We extract three different feature categories: sleeve length, neckline type, and item name, as illustrated in Fig. 9. Subsequently, we generate captions based on this feature information, for example, “short sleeve off shoulder t-shirts”.
Figure 10: Comparison between IDM–VTON and OOTDiffusion [52] on VITON-HD and DressCode dataset. Both methods are trained on VITON-HD training data. Best viewed in zoomed, color monitor.
Appendix 0.B Comparison with Concurrent Work
Figure 11: Comparison between IDM–VTON and OOTDiffusion [52] on In-the-Wild dataset. We provide visual comparison results on In-the-Wild dataset. Best viewed in zoomed, color monitor.
Table 5: Quantitative results on VITON-HD and DressCode dataset. We additionally compare the metric scores of IDM–VTON (ours) with the concurrent work OOTDiffusion [52].
Dataset VITON-HD DressCode
OOTDiffusion [52] 0.120 0.846 6.39 0.858 0.087 0.901 10.68 0.852
IDM–VTON (ours) 0.102 0.870 6.29 0.883 0.062 0.920 8.64 0.904
Table 6: Quantitative results on In-the-Wild dataset. We compare IDM–VTON (ours) with the concurrent work OOTDiffusion [52] on In-the-Wild dataset to assess the generalization capabilities. We report LPIPS, SSIM and CLIP image similarity scores.
Method LPIPS $\downarrow$ SSIM $\uparrow$ CLIP-I $\uparrow$
OOTDiffusion [52] 0.203 0.752 0.861
IDM–VTON (ours) 0.164 0.795 0.901
We additionally compare IDM–VTON with OOTDiffusion [52], which is a concurrent work on virtual try-on task. Both methods utilize an additional UNet to extract features from garment images. However, remark that there are differences between OOTDiffusion and IDM–VTON. Specifically, while OOTDiffusion trains the entire set of parameters of the additional UNet for garment feature extraction (i.e., corresponds to GarmentNet of IDM–VTON), IDM–VTON does not update GarmentNet. We find that pretrained UNet suffices in extracting low-level features, thus our approach is more efficient in terms of cost and memory.
For comparison with OOTDiffusion, we generate try-on images of OOTDiffusion at a resolution of $1024\times 768$, utilizing the model checkpoint trained on VITON-HD dataset, which is available via the official repository [^4] https://github.com/levihsu/OOTDiffusion. Fig. 10 and Fig. 11 present qualitative comparisons of OOTDiffusion and ours on VITON-HD, DressCode and In-the-Wild dataset. As shown in Fig. 10, We see that IDM–VTON outperforms OOTDiffusion in capturing both high-level semantics and low-level details, and generating more authentic images. In particular, we observe notable improvements of IDM–VTON compared to OOTDiffusion on In-the-Wild dataset, which demonstrates the generalization ability of IDM–VTON.
Tab. 5 and Tab. 6 show quantitative comparisons between IDM–VTON and OOTDiffusion on VITON-HD, DressCode, and In-the-Wild datasets. One can notify that IDM–VTON outperforms OOTDiffusion on all metrics including image fidelity (FID), and reconstruction of garments (LPIPS, SSIM and CLIP-I), which verifies our claim.
Figure 12: Qualitative comparison on VITON-HD and DressCode dataset. As we observed in our quantitative analysis, GAN-based methods generally struggle to generate high-fidelity images introducing non desirable distortions (e.g., non-realistic body and arms) while diffusion-based methods fail to capture low-level features or high-level semantics. All methods are trained on VITON-HD training data.
Figure 13: Qualitative comparison on In-the-Wild dataset. While baselines fail to generate natural images or capture the details of clothing, IDM–VTON produces realistic images and preserves fine details effectively. All methods are trained on VITON-HD training data.
Figure 14: Try-on results on VITON-HD test data by IDM–VTON trained on VITON-HD training data. Best viewed in zoomed, color monitor.
Figure 15: Try-on results on DressCode test data by IDM–VTON trained on VITON-HD training data. Best viewed in zoomed, color monitor.
Figure 16: Try-on results on In-the-Wild dataset by IDM–VTON trained on VITON-HD training data. Best viewed in zoomed, color monitor.
Figure 17: Try-on results on In-the-Wild dataset by IDM–VTON trained on VITON-HD training data. Best viewed in zoomed, color monitor.
Figure 18: Try-on results on In-the-Wild dataset by IDM–VTON trained on VITON-HD training data. Best viewed in zoomed, color monitor.